OpenStudy (anonymous):

Whats the volume of the solid cutted from the sphere defined by x^2 + y^2 + z^2=4 by the cilinder defined by r=2sin theta ?

OpenStudy (anonymous):

\[\int\limits_{0}^{\Pi}\int\limits_{0}^{2\sin \theta}\int\limits_{-\sqrt{4-r^2}}^{\sqrt{4-r^2}} r dzdrd \Theta\]

OpenStudy (anonymous):

do i got this right?

OpenStudy (anonymous):

Ages since I did this stuff: How about 4 int (-4 to 2sin theta) int (0 to 4-x^2) sqrt(4-x^2-y^2) dxdy. (maybe).

OpenStudy (anonymous):

I guess thats valid too, but i think it might be a shortcut to do this with cilindric coords... im gonna try both anyways. Thank you!